OpenStudy (anonymous):

will anyone be kind and help me with two problems ;) thanks if you will lovelys<3

jimthompson5910 (jim_thompson5910):

\[\Large 3\sqrt[5]{(x+2)^3}+3 = 27\] \[\Large 3\sqrt[5]{(x+2)^3} = 27-3\] \[\Large 3\sqrt[5]{(x+2)^3} = 24\] \[\Large \sqrt[5]{(x+2)^3} = \frac{24}{3}\] \[\Large \sqrt[5]{(x+2)^3} = 8\] \[\Large (x+2)^3 = 8^5\] \[\Large (x+2)^3 = 32768\] \[\Large x+2 = \sqrt[3]{32768}\] \[\Large x+2 = 32\] \[\Large x = 32-2\] \[\Large x = 30\]

OpenStudy (anonymous):

thanks a lot jim!

jimthompson5910 (jim_thompson5910):

np